OpenStudy (anonymous):

Bill wants to factor x^2+8x+16 by grouping; however sara says it is a special product and can factor a different way. using complete sentences, explain and demonstrate how both methods will result in the same factors.

OpenStudy (anonymous):

It's a perfect square, I would just factor by inspection. That could be one method, lol

OpenStudy (anonymous):

I suck at math so that makes no sense.

OpenStudy (zzr0ck3r):

(a+b)^2=a^2+2ab+b^2

OpenStudy (anonymous):

Hm. So you know that since there's no coefficient before x^2 your factor starts with (x and the number(s) following need to multiply by to give you 16 and add to give you 8--and 4 work, so it's a perfect square.

OpenStudy (anonymous):

okie so what does it mean by sara saying it's a special product?

OpenStudy (anonymous):

It's special because your factors are the same. :)

OpenStudy (anonymous):

So what am I suppose to say??

OpenStudy (anonymous):

I don't understand what this question wants me to do!!!

OpenStudy (anonymous):

Please somebody help me???

zepdrix (zepdrix):

So they want you to use two different methods. Let's look at Bill's method first (grouping).\[\Large\bf\sf x^2+8x+16\quad=\quad x^2+4x+4x+16\]8x can be written as 4x and 4x. From here we'll factor an x out of each of the first two terms,\[\Large\bf\sf x(x+4)+4x+16\]And factor a 4 out of each of the last two terms,\[\Large\bf\sf x(x+4)+4(x+4)\]

zepdrix (zepdrix):

From there we can see that they each have a factor of (x+4), so we'll factor that out of each, \[\Large\bf\sf \color{orangered}{x}(x+4)\color{orangered}{+4}(x+4)\quad=\quad (x+4)(\color{orangered}{x+4})\]

zepdrix (zepdrix):

The same factor is being multiplied by itself, so we can write it as a square,\[\Large\bf\sf (x+4)(x+4)\quad=\quad (x+4)^2\]

zepdrix (zepdrix):

So that's Bill's method. Make sensssseeeee, kinda?

OpenStudy (anonymous):

Yeah. kind of. math is like learning a very old language to me at times. So what would sara's method be?

zepdrix (zepdrix):

Sara remembers that there is a special way to simplify the `sum of squares`. \[\Large\bf\sf a^2+2ab+b^2\quad=\quad (a+b)^2\]So we start with:\[\Large\bf\sf x^2+8x+16\]We need to write 16 as a square, and also need to write 8x as `2 times something`. And hopefully that something will be our a and b.

zepdrix (zepdrix):

16 is 4^2 right? And 8x is 2*4x.

zepdrix (zepdrix):

\[\Large\bf\sf x^2+8x+16\quad=\quad x^2+2(4x)+4^2\]

zepdrix (zepdrix):

See how we can use the rule for the `sum of squares` to simplify it?

zepdrix (zepdrix):

\[\Large\bf\sf a^2+2(ab)+b^2\quad=\quad (a+b)^2\]\[\Large\bf\sf x^2+2(4x)+4^2\quad = \quad ?\]

OpenStudy (anonymous):

Sorry... but I am totally lost. In math... I feel like I should be in 5th grade instead of 12th!!

zepdrix (zepdrix):

:c

OpenStudy (anonymous):

srry

zepdrix (zepdrix):

This is the rule for the sum of squares:\[\Large\bf\sf \color{royalblue}{a}^2+2(\color{royalblue}{a}\color{orangered}{b})+\color{orangered}{b}^2\quad=\quad (\color{royalblue}{a}+\color{orangered}{b})^2\] Understand how we can apply it to our problem here? \[\Large\bf\sf \color{royalblue}{x}^2+2(\color{orangered}{4}\color{royalblue}{x})+\color{orangered}{4}^2\quad=\quad ?\] We're trying to turn it into something like the rule shows on the right.

OpenStudy (anonymous):

so it's (x+4)^2

zepdrix (zepdrix):

Yes, good.

OpenStudy (anonymous):

so what do I do next?

zepdrix (zepdrix):

Both Bill and Sara's methods resulted in (x+4)^2, the same answer. Explain in sentences how you got the same answer using different methods. Use my steps to try and explain it.

OpenStudy (anonymous):

How is this??? Bill got (x+4)^2 by doing x^2+4x+4x+16. He than factored out an x from each of the first two terms, getting x(x+4) + 4x+16. Next he factored out a 4 out of each of the last two terms, x(x+4) + 4(x+4). He than factored out (x+4), since they each have it. Bill then got (x+4)(x+4), which he then squared it being that the same factor was being multiplied by itself. He ends the problem with (x+4)^2. Sara knows that a^2+2ab+b^2= (a+b)^2. So Sara starts with x^2+8x+16, and knows that she needs to write 16 as a square and 8x will be 2 times 4x. So Sara gets, x^2+2(4x)+4^2. Being that the rule for sum of squares is a^2+2(ab)+b^2=(a+b)^2, Sara knows that the x in X^2 is the a and the 4s in 2(4x)+4^2 is the b's, she gets (x+4)^2.

zepdrix (zepdrix):

lol yayyyy good job! \c:/

OpenStudy (anonymous):

So that makes sense??

zepdrix (zepdrix):

Just seems like you copied what I wrote lolol XD Yes it makes sense, I'm hoping it makes at least a little sense to you :3

OpenStudy (anonymous):

Yeah going throught it helped.

OpenStudy (anonymous):

Can you help me with this question? Three functions are given below: f(x), g(x), and h(x). Explain how to find the axis of symmetry for each function and rank the functions based on their axis of symmetry (from smallest to largest). (10 points) f(x) = 3(x + 4)2+ 1 g(x) = 2x2 -16x + 15 h(x) is graphed.

zepdrix (zepdrix):

So if you look at the graph, where is the line of symmetry for the curve drawn? In other words: Where could we fold the paper in half to make the curve overlap itself perfectly.

OpenStudy (anonymous):

(1,-3)

zepdrix (zepdrix):

Ok good that's where the vertex of the parabola is located. But if want an axis of symmetry then we need a line that we can fold over. So we want that entire line going up, x=1. See how the graph is mirrored on both sides of that line?

OpenStudy (anonymous):

yeah.

zepdrix (zepdrix):

Notice how the axis of symmetry was simply the `x coordinate of our vertex`? The other functions will work out just as easily. Do you know how to find the vertex of f(x) and g(x)?

OpenStudy (anonymous):

I understand this question thanks. I have had the following question for a while... Sandra exclaims that her quadratic with a discriminant of -4 has no real solutions. Sandra then puts down her pencil and refuses to do any more work. Create an equation with a negative discriminant. Then explain to Sandra, in calm and complete sentences, how to find the solutions, even though they are not real. (10 points)

zepdrix (zepdrix):

Woah calm down Sandra! Let's find the complex solutions as well! :U

zepdrix (zepdrix):

So we need to start with an equation that has a negative discriminant? Umm sec lemme think of something.

zepdrix (zepdrix):

\[\Large\bf\sf x^2-2x+5\] If we throw this polynomial into the Quadratic Formula:\[\Large\bf\sf x\quad=\quad\frac{-b\pm \sqrt{b^2-4ac}}{2a}\quad=\quad \frac{2\pm\sqrt{4-4\cdot1\cdot5}}{2(1)}\]

zepdrix (zepdrix):

Our discriminant is the stuff under the radical. \(\Large\bf\sf 4-4\cdot5\quad=\quad 4-20\quad=\quad -16\) Negative is bad when under root, yes?

OpenStudy (anonymous):

um.... yes?

zepdrix (zepdrix):

Yes. We can't take the square root of a negative number. Bad. But we have a way of classifying these bad numbers by calling them "imaginary". The way we do this is by calling \(\Large\bf\sf \sqrt{-1}=\mathcal i\) .

zepdrix (zepdrix):

Here is a way we can deal with the negative under the square root,\[\Large\bf\sf \sqrt{-16}\quad=\quad \sqrt{-1\cdot 16}\quad=\quad \sqrt{-1}\cdot \sqrt{16}\]See how we were able to separate the bad number from the rest of our stuff?

zepdrix (zepdrix):

\[\Large\bf\sf \sqrt{-1}\cdot\sqrt{16}\quad=\quad \mathcal i\cdot \sqrt{16}\]

zepdrix (zepdrix):

Applying the Quadratic Formula to our quadratic function we get,\[\Large\bf\sf x\quad=\quad \frac{2\pm\sqrt{16}\cdot \mathcal i}{2}\]Which can be simplified down a bit. I think 16 is a perfect square isn't it? Can we take it out of the root somehow?

OpenStudy (anonymous):

I don't think so??

zepdrix (zepdrix):

lol :)

OpenStudy (anonymous):

wat?

zepdrix (zepdrix):

\[\Large\bf\sf 4\cdot 4\quad=\quad 16\] \[\Large\bf\sf \implies \quad \sqrt{16}\quad=\quad 4\]

zepdrix (zepdrix):

Square roots too confusing? :o

OpenStudy (anonymous):

no. well yeah kinda. I am just trying to type in my responses to the first question and another. so the square root of 16 is 4...

zepdrix (zepdrix):

Yes. So our expression further simplifies.\[\Large\bf\sf x\quad=\quad \frac{2\pm\sqrt{16}\cdot \mathcal i}{2}\qquad\to\qquad x\quad=\quad \frac{2\pm 4 \mathcal i}{2}\]Which can be further simplified. Break it into two separate fractions,\[\Large\bf\sf x\quad=\quad \frac{2}{2}\pm\frac{4\mathcal i}{2}\]

zepdrix (zepdrix):

Do you understand how to further simplify the fractions?

OpenStudy (anonymous):

well if somebody asked me to show them how I would say I don't know but I am guessing the square root 16*I simplfys down to 4i and so on.

OpenStudy (anonymous):

ya still there?

OpenStudy (anonymous):

zepdrix???? I still need help!!!